Find the range space and the kernel of the
linear transformation :
T: R4 --> R4, T(x1, x2, x3, x4) =
(x1+ x2+ x3+ x4, x1+ x2, x3+ x4, 0)
1
Expert's answer
2020-03-31T10:29:28-0400

"T: \R^4\to \R^4,"

"T(x_1,x_2,x_3,x_4)=(x_1+x_2+x_3+x_4,x_1+x_2,x_3+x_4,0)." (1)


For "T\in L(\R^4,\R^4)," the range of "T" is the subset of "\R^4" consisting of those vectors that are of the form "T(x_1,x_2,x_3,x_4)" for some "(x_1,x_2,x_3,x_4)\in \R^4:"


"range T=\{ T(x_1,x_2,x_3,x_4): (x_1,x_2,x_3,x_4)\in \R^4\}."


So, "T\in L(\R^4,\R^4)" is defined by (1), then

"range T=\{(x_1+x_2+x_3+x_4,x_1+x_2,x_3+x_4,0): x_1,x_2,x_3,x_4\in\R\}."


"A=\begin{bmatrix}
 1 & 1 & 1 & 1 \\
 1 & 1 & 0 & 0 \\
 0 & 0 & 1 & 1\\ 
 0 & 0 & 0 & 0 
\end{bmatrix}"

is  the matrix satisfying "T(x)=Ax, x\in\R^4." The range of the linear transformation "T" is the same as the range of the matrix "A."

The range of "A" is the columns space of "A" . Thus it is spanned by columns

"\begin{bmatrix}
 1 \\
 1 \\
 0 \\
 0 
\end{bmatrix}, \begin{bmatrix}
 1 \\
 0 \\
 1 \\
 0 
\end{bmatrix}"

From the below (2) reduction of the augmented matrix, we see that these vectors are linearly independent, thus a basis for the range.

Therefore, we have

"range T= range A=span\{\begin{bmatrix}
 1 \\
 1 \\
 0 \\
 0 
\end{bmatrix},\begin{bmatrix}
 1 \\
 0 \\
 1 \\
 0 
\end{bmatrix}\}"

and

"\{\begin{bmatrix}
 1 \\
 1 \\
 0 \\
 0 
\end{bmatrix},
\begin{bmatrix}
 1 \\
 0 \\
 1 \\
 0 
\end{bmatrix}\}"

is a basis for "range T."

"\ker T=\{(x_1,x_2,x_3,x_4)\in\R^4: T(x_1,x_2,x_3,x_4)=0\}."

System of LE associated to the implicit equations of the kernel, resulting from equalling to zero the components of the linear transformation formula.

"x_1+x_2+x_3+x_4=0"

"x_1+x_2=0"

"x_3+x_4=0"

Transform the augmented matrix "A" to row echelon form.

"r_2\to r_2-r_1"

"\begin{bmatrix}
 1 & 1 & 1 & 1 \\
 0 & 0 & -1 & -1 \\
 0 & 0 & 1 & 1\\ 
 0 & 0 & 0 & 0 
\end{bmatrix}"

"r_3\to r_3+r_2"

"\begin{bmatrix}
 1 & 1 & 1 & 1 \\
 0 & 0 & -1 & -1 \\
 0 & 0 & 0 & 0\\ 
 0 & 0 & 0 & 0 
\end{bmatrix}"

"r_1\to r_1+r_2"

"\begin{bmatrix}
 1 & 1 & 0 & 0 \\
 0 & 0 & -1 & -1 \\
 0 & 0 & 0 & 0\\ 
 0 & 0 & 0 & 0 
\end{bmatrix}"

"r_2\to -1r_2"

"\begin{bmatrix}
 1 & 1 & 0 & 0 \\
 0 & 0 & 1 & 1 \\
 0 & 0 & 0 & 0\\ 
 0 & 0 & 0 & 0 
\end{bmatrix}" (2)

Translate the row echelon form matrix to the associated system of linear equations

"x_1+x_2=0"

"x_3+x_4=0"

The implicit equations of the kernel are the equations above


"\ker T=\{(x_1,x_2,x_3,x_4)\in\R^4 | x_1+x_2=0; x_3+x_4=0 \}"


"\ker T =\begin{bmatrix}
 s \\
 -s \\
t \\
-t 
\end{bmatrix}: t,s\in\R,"

"\ker T=span\{\begin{bmatrix}
 1 \\
 -1 \\
0 \\
0
\end{bmatrix},
\begin{bmatrix}
 0 \\
 0 \\
1 \\
-1
\end{bmatrix}\}"

and

"\{\begin{bmatrix}
 1 \\
 -1 \\
 0 \\
 0 
\end{bmatrix},
\begin{bmatrix}
 0 \\
 0 \\
 1 \\
 -1 
\end{bmatrix}\}" is a basis for "\ker T."